# Verifying Equilibria in Finite-Horizon Probabilistic Concurrent Game Systems

### v4

@@ 1. Introduction
The ubiquity of distributed systems in modern computing architectures has motivated recent work into the study of multi-agent systems [GTW02, Gup16, MMPV14, SL09, VdHW08, Woo09, AKM+19, RBV23, RV22, RV21, BBM10, BBMU11, BMS14, BBMU15, BRRvdB24, BRvdB23], which are abstract mathematical models in which two or more rational agents interact with one another in order to achieve their goals. The broad motivation behind these works is to provide a mathematical framework for analyzing the overall trends of the complex mechanics that arise when multiple agents interact with one another in a strategic manner.
Much of the relevant work in this area comes from the field of formal methods, which has recently begun to analyze equilibria in multi-agent systems as a way to analyze “stable” behaviors in multi-agent systems, and, therefore, analyze the system’s properties as a whole. The two major problems in the field of formal methods are the realizability [PR89] and the verification problems [Ros92, RV22]. Therefore, much of the work in formal analysis of multi-agent systems is concerned with determining the existence of (the realizability problem) and verifying (the verification problem) different types of equilibria [GTW02, BBMU11, BBM10, BBMU15, RBV23, RV22, RV21, Gup16, GNPW18, GPW17, BBMU15, BRRvdB24, BRvdB23]. However, as evidenced by the previously cited works, work in formal verification often comes with a strong preference for deterministic systems and deterministic agent behaviors. This creates something of an ideological divide between equilibrium analysis in the formal methods literature and the study of equilibria in the game theory literature. Our broad goal in this paper is to begin to bridge this divide by applying this formal methods style of equilibrium analysis to probabilistic systems. This is a new line of inquiry, as even papers in the formal methods literature that consider probabilistic systems such as [GHL+21] (which itself notes the focus on deterministic systems in previous literature) consider notions of almost-sure satisfaction on infinite-length executions of a system. These notions allow for these probabilistic systems to be analyzed through the same techniques used for deterministic ones.
@@
In order to accurately isolate the complexity of verification, we introduce a new succinctly represented model in this paper, the b-bounded concurrent concurrent game system. In a b-bounded system, only b agents select an action concurrently at all given game-states, where b is a fixed constant. This model serves two purposes. First, it allows us to generalize from the popular but restrictive turn-based game model that is ubiquitous in the literature [Umm11, Gup16, Woo09, Rou10]. Second, it sharpens our complexity-theoretic results by circumventing the construction of an exponentially large transition table as part of our input. In a multi-agent system, the transition table of an unrestricted model can be seen as an exponential construction in itself, a point we discuss explicitly in Section 2 and that has been noted in other works such as [BBMU15]. This exponential construction within the input devalues the complexity-theoretic value of upper bounds, as a PSPACE upper bound would practically translate into doubly-exponential time complexity in the number of agents. We discuss how our results extend to a full concurrency model in Section 3.6 (one in which all agents select an action at all game-states). Furthermore, the issue of goal representation also plays a significant role in the overall computation complexity of the verification problem. As noted in [RV21, RV22], if agents are given highly succinct temporal logic formulae that represent their preferences on the system’s executions, then the worst-case complexity of reasoning about these logics can completely overpower the worst-case complexity of the actual verification decision problem. In order to avoid this problem, we give the agents reachability goals, which are arguably the simplest non-trivial goals.
@@ 2. Preliminaries
    $V$ is a set of states, with $v_{0}\in V$ the initial state.
@@
    $A=\{A_{1},A_{2},\ldots A_{k}\}$ is the set of action sets, where Agent $i$ is associated with the set of actions $A_{i}$. We refer to cross-product spaces $\Theta_{W}=\bigtimes_{i\in W}A_{i}$ for some $W\subseteq\Omega$ and then use $\theta$ to refer to an element of $\Theta_{W}$.
@@
    $P:V\rightarrow 2^{\Omega}$ is the playing function. This function decides which agents are “active” at a given state, as not every agent makes a relevant decision at every state. The game being $b$-bounded means that at most $b$ agents can make a relevant decision at all states, i.e., $|P(v)|\leq b$ for all $v\in V$ for $b\geq 1$
@@
Note that in this definition, we have $1$-indexed the set of agents as opposed to 00-indexing that is more present in the related computer science literature. This is because the model presented more closely lines up with models studied in game theory [Sol22, FT91], and the game theory literature adopts mathematical conventions such as $1$-indexing as opposed to computer science conventions such as 00-indexing.
@@
The game starts at the state $v_{0}$. At this state, the agents in $P(v_{0})$ choose a distribution over the actions from their action set. This collective choice is then input into the transition function $\delta$, which returns a distribution over a set of new states. This distribution is then sampled, and the execution of the system proceeds to the next state $v^{\prime}$, where the process repeats with the agents in $P(v^{\prime})$ until $F$ states have been visited (it is acceptable to visit a single state multiple times). After the $F-1$-th action choice, the game terminates, and no further moves are made. The entire record of the $F-1$ action choices and the corresponding states visited is called a play.
Agents are incentivized to choose their actions so that a state in their reachability goal is eventually seen. In other words, agents prefer plays that visit a state in their reachability goal to those that do not. For an Agent $i$, a play that visits a state in its reachability set yields a payoff of $1$, whereas a play that does not visit a state in its reachability set yields a payoff of 00. Agents select their actions to maximize their expected payoffs.
@@
    The set of states $S$ is given by a cross product of the component state-sets, $S=\bigtimes_{i\in\Omega}S_{i}$. $s_{0}\in S$ is given by $\langle s^{0}_{0},\ldots,s^{k-1}_{0}\rangle$
@@
    The output function $O$ aggregates the component-wise output functions. The output $O(\langle s^{1},\ldots,s^{k}\rangle,v)$ is given by $,\langle O_{1}(s^{1},v),$$\ldots,O_{k}(s^{k},v)\rangle$. For $i\in P(v)$, Each component $O_{i}(s^{i},v)$ belongs to $\mathbb{D}(A_{i})$. For $j\not\in P(v)$, the output $O_{i}(s^{j},v)$ is $\bot$. By a slight abuse of notation, we project these $\bot$ values out, and so, taken together, $O(s,v)\in\bigtimes_{i\in P(v)}\mathbb{D}(A_{i})$.
@@
The payoff for Agent $i$ under a strategy profile $\pi$ denoted $p(\pi,i)$ can be calculated $p(\pi,i)=\sum_{t\in\mathrm{Plays}}\mathbb{P}(\pi,t)*\mathrm{Payoff}(t,i)$ where $\mathbb{P}(\pi,t)$ refers to the probability that when all agents follow $\pi$ the play $t$ is sampled and $\mathrm{Payoff}(t,i)$ is the payoff Agent $i$ receives from the play $t$ ; 1 if it contains a state in $G_{i}$ and 00 otherwise. Given this notion of a payoff under a strategy profile, we can now define our first solution concept of interest.
@@
In other words, a Nash equilibrium is a strategy profile such that any unilateral change in strategy for any Agent $i$, from $\pi_{i}$ to some other $\pi^{\prime}_{i}$, does not result in a greater probability that Agent $i$’s goal is satisfied. Therefore, Agent $i$ is not incentivized to change her strategy unilaterally. The Nash equilibrium verification problem takes a concurrent game $\mathbb{G}$ and a strategy profile $\pi$ specified by a tuple of probabilistic transducers $\langle\pi_{1}\ldots\pi_{i}\ldots\pi_{k}\rangle$ as inputs, and the problem is to decide whether $\pi$ is a Nash equilibrium in $\mathbb{G}$.
@@
    For each pair of consecutive elements $h[j],h[j+1]$ in $h$, it must be the case that there exists some tuple of actions that transitions from $h[j]$ to $h[j+1]$ with positive probability, i.e., $\exists\theta\in\bigtimes_{i\in P(h[j])}A_{i}.\delta(h[j],\theta,h[j+1])>0$
@@
Given a concurrent game $\mathbb{G}$ and a history $h$. the subgame $\mathbb{G}|_{h}$ is the game $\mathbb{G}$ after having observed the history $h$. Its initial state is the last element of $h$, and depending on the contents of $h$, some agents might have a fixed payoff of $1$ in this subgame if $h$ visited one of their goal states. Furthermore, the finite horizon of $\mathbb{G}|_{h}$ is updated to $F-|h|$, where $F$ is the original finite horizon of $\mathbb{G}$.
@@
A strategy profile $\pi=\langle\pi_{1},\pi_{2},\ldots\pi_{k}\rangle$ is a subgame perfect equilibrium in a concurrent game $\mathbb{G}$ iff it is the case that for all histories $h$ we have that $\pi|_{h}:=\langle\pi_{1}|_{h},\pi_{2}|_{h},\ldots\pi_{k}|_{h}\rangle$ is a Nash equilibrium in $\mathbb{G}|_{h}$
@@ 3. Subgame Perfect Equilibrium Upper Bound
where $supp(\cdot)$ refers to the support of a probabilistic distribution, i.e., the set of elements that are assigned a non-zero probability by the distribution. The expression $\mathbb{P}(O(s_{1},v_{1})[\theta])$ represents the probability that, when sampled, $O(s_{1},v_{1})$ returns $\theta$. In our bounded-concurrency setting, this sum iterates over ×$\bigtimes_{i\in P(v_{1})}A_{i}$ – details about the full-concurrency setting can be found in Section 3.6. This set has polynomial size w.r.t the input due to the bound $b$, so it can be iterated through in PSPACE.
@@ 3.1. Outline
The overarching goal of this section is to prove that the problem of deciding whether a strategy profile $\pi$ is a subgame perfect equilibrium in a $b$-bounded concurrent game $\mathbb{G}$ is in PSPACE. In order to accomplish this, we first recall that PSPACE is closed under non-determinism and complementation, i.e. PSPACE=NPSPACE=co-PSPACE. Since PSPACE is closed under complementation, we consider the problem of determining whether $\pi$ is not a subgame perfect equilibrium in $\mathbb{G}$.
@@ 3.2. Histories and Subgames
Starting from the state $\langle v_{0},s_{0},1\rangle$, a path from $\langle v_{0},s_{0},1\rangle$ to $\langle v_{r},s_{r},n_{r}\rangle$ can be guessed state by state on-the-fly in NPSPACE-PSPACE. At each state $\langle v_{j},s_{j},n_{j}\rangle$, the algorithm can guess a successor $\langle v_{j+1},s_{j+1},n_{j+1}\rangle$ and verify that $\gamma(s_{j},v_{j})=s_{j+1}$, $n_{j+1}=n_{j}+1$ and there exists a global action $\theta\in\bigtimes_{k\in P(v_{j})}A_{k}$ such that $\delta(v_{j},\theta,v_{j+1})>0$ in PSPACE by iterating over all possibilities for $\theta$. It can also verify that $v_{j}\not\in G_{i}$. The algorithm only needs to store at most two states in its memory – the current state and the guessed successor. Since the state space $V\times S\times[F]$ is exponential in the size of the input, individual states can be encoded in PSPACE.
The set $R$ of relevant history reachable states represents the states that occur in $\mathbb{G}\times\pi$ after some history $h$ in $\mathbb{G}$ has been observed. Therefore, if we want to analyze the payoff given to Agent $i$ in a subgame induced by a history $h$ in $\mathbb{G}$, we consider $\mathbb{G}\times\pi$ started from the state $r\in R$ that results from following the history $h$.
@@ 3.3. Computing Winning Probabilities From a State
Consider a Markov Chain $\mathcal{C}$ with two non-target states $x$ and $y$ with the following properties : The state $x$ self transitions with probability $\frac{1}{2}$ and transitions to $y$ with probability $\frac{1}{2}$. Furthermore, the state $y$ self transitions with probability $\frac{1}{2}$ and transitions to $x$ with probability $\frac{1}{2}$. The equations for the hitting probabilities $h_{x}$ and $h_{y}$ would be both given by the linear equation $\frac{1}{2}h_{x}+\frac{1}{2}h_{y}$, meaning that the system of equations used to solve for hitting probabilities in $\mathcal{C}$ would be underdetermined and have infinite solutions.
However, the Markov Chain $\mathbb{G}\times\pi$ has a special structure. Since it explicitly tracks the time index $n\in[F]$, this means that it contains no cycles – states with time index $n$ can only transition to states that have time index $n+1$.
@@
We demonstrate this by backwards induction. Consider the set of states $\langle v,s,F\rangle$, i.e. the set of states with time index $n=F$. These states have no outgoing transitions, so we have $h_{\langle v,s,F\rangle}=1$ if $v\in G_{i}$ and 00 otherwise. Therefore, all states of the form $\langle v,s,F\rangle$ have unique solutions for their hitting probabilities.
@@
[…]

### v5

@@ 1. Introduction
The ubiquity of distributed systems in modern computing architectures has motivated much recent work into the study of multi-agent systems [GTW02, Gup16, MMPV14, SL09, VdHW08, Woo09, AKM+19, RBV23, RV22, RV21, BBM10, BBMU11, BMS14, BBMU15, BRRvdB24, BRvdB23], which are abstract mathematical models in which two or more rational agents interact with each other in order to achieve their goals. The broad motivation behind these works is to provide a mathematical framework for analyzing the overall trends of the complex mechanics that arise when multiple agents interact with one another in a strategic manner.
Much of the relevant work in this area comes from the field of formal methods, which has recently begun to analyze equilibria in multi-agent systems as a way to analyze “stable” behaviors in multi-agent systems, and, therefore, analyze the system’s properties as a whole. The two major problems in the field of formal methods are the realizability [PR89] and the verification problems [Ros92, RV22]. Therefore, much of the work in formal analysis of multi-agent systems is concerned with determining the existence of (the realizability problem) and verifying (the verification problem) different types of equilibria [GTW02, BBMU11, BBM10, BBMU15, RBV23, RV22, RV21, Gup16, GNPW18, GPW17, BBMU15, BRRvdB24, BRvdB23]. As evidenced by the previously cited works, however, work in formal verification often comes with a strong preference for deterministic systems and deterministic agent behaviors. This creates something of an ideological divide between equilibrium analysis in the formal methods literature and the study of equilibria in the game theory literature. Our broad goal in this paper is to begin to bridge this divide by applying this formal methods style of equilibrium analysis to probabilistic systems. This is a new line of inquiry, as even papers in the formal methods literature that consider probabilistic systems such as [GHL+21] (which itself notes the focus on deterministic systems in previous literature) consider notions of almost-sure satisfaction on infinite-length executions of a system. These notions allow for these probabilistic systems to be analyzed through the same techniques used for deterministic ones.
@@
In order to accurately isolate the complexity of verification, we introduce a new succinctly represented model in this paper, the b-bounded concurrent game system. In a b-bounded system, only b agents select an action concurrently at all given game-states, where b is a fixed constant. This model serves two purposes. First, it allows us to generalize from the popular but restrictive turn-based game model that is ubiquitous in the literature [Umm11, Gup16, Woo09, Rou10]. Second, it sharpens our complexity-theoretic results by circumventing the construction of an exponentially large transition table as part of our input. In a multi-agent system, the transition table of an unrestricted model can be seen as an exponential construction in itself, a point we discuss explicitly in Section 2 and that has been noted in other works such as [BBMU15]. This exponential construction within the input devalues the complexity-theoretic value of upper bounds, as a PSPACE upper bound would practically translate into doubly-exponential time complexity in the number of agents. We discuss how our results extend to a full concurrency model in Section 3.6 (one in which all agents select an action at all game-states). Furthermore, the issue of goal representation also plays a significant role in the overall computation complexity of the verification problem. As noted in [RV21, RV22], if agents are given highly succinct temporal logic formulae that represent their preferences on the system’s executions, then the worst-case complexity of reasoning about these logics can completely overpower the worst-case complexity of the actual verification decision problem. In order to avoid this problem, we give the agents reachability goals, which are arguably the simplest non-trivial goals.
@@ 2. Preliminaries
    $V$ is a finite set of states, with $v_{0}\in V$ the initial state.
@@
    $A=\{A_{1},A_{2},\ldots A_{k}\}$ is the set of finite action sets, where Agent $i$ is associated with the set of actions $A_{i}$. We refer to cross-product spaces $\Theta_{W}=\bigtimes_{i\in W}A_{i}$ for some $W\subseteq\Omega$ and then use $\theta$ to refer to an element of $\Theta_{W}$.
@@
    $P:V\rightarrow 2^{\Omega}$ is the playing function. This function decides which agents are “active” at a given state, as not every agent makes a relevant decision at every state. The game being $b$-bounded means that at most $b$ agents can make a relevant decision at all states, i.e., $|P(v)|\leq b$ for all $v\in V$ for $b\geq 1$.
@@
Note that in this definition, we have $1$-indexed the set of agents as opposed to 0-indexing that is more present in the related computer science literature. This is because the model presented more closely lines up with models studied in game theory [Sol22, FT91], and the game theory literature adopts mathematical conventions such as $1$-indexing as opposed to computer science conventions such as 0-indexing.
@@
The game starts at the state $v_{0}$. At this state, the agents in $P(v_{0})$ choose a distribution over the actions from their action set. This collective choice is then input into the transition function $\delta$, which returns a distribution over a set of new states. This distribution is then sampled, and the execution of the system proceeds to the next state $v^{\prime}$, where the process repeats with the agents in $P(v^{\prime})$ until $F$ states have been visited (it is acceptable to visit a single state multiple times). After the F1F-1-th action choice, the game terminates, and no further moves are made. The entire record of the F1F-1 action choices and the corresponding states visited is called a play.
Agents are incentivized to choose their actions so that a state in their reachability goal is eventually seen. In other words, agents prefer plays that visit a state in their reachability goal to those that do not. For an Agent $i$, a play that visits a state in its reachability set yields a payoff of $1$, whereas a play that does not visit a state in its reachability set yields a payoff of 0. Agents select their actions to maximize their expected payoffs.
@@
    The set of states $S$ is given by a cross product of the component state-sets, $S=\bigtimes_{i\in\Omega}S_{i}$. $s_{0}\in S$ is given by $\langle s^{0}_{0},\ldots,s^{k-1}_{0}\rangle$.
@@
    The output function $O$ aggregates the component-wise output functions. The output $O(\langle s^{1},\ldots,s^{k}\rangle,v)$ is given by $\langle O_{1}(s^{1},v),$ $\ldots,O_{k}(s^{k},v)\rangle$. For $i\in P(v)$, Each component $O_{i}(s^{i},v)$ belongs to $\mathbb{D}(A_{i})$. For $j\not\in P(v)$, the output $O_{i}(s^{j},v)$ is $\bot$. By a slight abuse of notation, we project these $\bot$ values out, and so, taken together, $O(s,v)\in\bigtimes_{i\in P(v)}\mathbb{D}(A_{i})$.
@@
The payoff for Agent $i$ under a strategy profile $\pi$ denoted $p(\pi,i)$ can be calculated $p(\pi,i)=\sum_{t\in\mathrm{Plays}}\mathbb{P}(\pi,t)*\mathrm{Payoff}(t,i)$ where $\mathbb{P}(\pi,t)$ refers to the probability that when all agents follow $\pi$ the play $t$ is sampled and $\mathrm{Payoff}(t,i)$ is the payoff Agent $i$ receives from the play $t$ ; 1 if it contains a state in $G_{i}$ and 0 otherwise. Given this notion of a payoff under a strategy profile, we can now define our first solution concept of interest.
@@
In other words, a Nash equilibrium is a strategy profile such that any unilateral change in strategy for any Agent $i$, from $\pi_{i}$ to some other $\pi^{\prime}_{i}$, does not result in a greater probability that Agent $i$’s goal is satisfied. Therefore, Agent $i$ is not incentivized to change her strategy unilaterally. The Nash-equilibrium verification problem takes a concurrent game $\mathbb{G}$ and a strategy profile $\pi$ specified by a tuple of probabilistic transducers $\langle\pi_{1}\ldots\pi_{i}\ldots\pi_{k}\rangle$ as inputs, and the problem is to decide whether $\pi$ is a Nash equilibrium in $\mathbb{G}$.
@@
    For each pair of consecutive elements $h[j],h[j+1]$ in $h$, it must be the case that there exists some tuple of actions that transitions from $h[j]$ to $h[j+1]$ with positive probability, i.e., $\exists\theta\in\bigtimes_{i\in P(h[j])}A_{i}.\delta(h[j],\theta,h[j+1])>0$.
@@
Given a concurrent game $\mathbb{G}$ and a history $h$. the subgame $\mathbb{G}|_{h}$ is the game $\mathbb{G}$ after having observed the history $h$. Its initial state is the last element of $h$, and depending on the contents of $h$, some agents might have a fixed payoff of $1$ in this subgame if $h$ visited one of their goal states. Furthermore, the finite horizon of $\mathbb{G}|_{h}$ is updated to F|h|F-|h|, where $F$ is the original finite horizon of $\mathbb{G}$.
@@
A strategy profile $\pi=\langle\pi_{1},\pi_{2},\ldots\pi_{k}\rangle$ is a subgame perfect equilibrium in a concurrent game $\mathbb{G}$ iff it is the case that for all histories $h$ we have that $\pi|_{h}:=\langle\pi_{1}|_{h},\pi_{2}|_{h},\ldots\pi_{k}|_{h}\rangle$ is a Nash equilibrium in $\mathbb{G}|_{h}$.
@@ 3. Subgame Perfect Equilibrium Upper Bound
where $supp(\cdot)$ refers to the support of a probabilistic distribution, i.e., the set of elements that are assigned a non-zero probability by the distribution. The expression $\mathbb{P}(O(s_{1},v_{1})[\theta])$ represents the probability that, when sampled, $O(s_{1},v_{1})$ returns $\theta$. In our bounded-concurrency setting, this sum iterates over $\bigtimes_{i\in P(v_{1})}A_{i}$ – details about the full-concurrency setting can be found in Section 3.6. This set has polynomial size w.r.t the input due to the bound $b$, so it can be iterated through in PSPACE.
@@ 3.1. Outline
The overarching goal of this section is to prove that the problem of deciding whether a strategy profile $\pi$ is a subgame perfect equilibrium in a $b$-bounded concurrent game $\mathbb{G}$ is in PSPACE. This consists of several steps, which are detailed in the following sections, and a broad overview of these sections and their goals can be found in Figure 1. In order to establish our PSPACE upper bound, we first recall that PSPACE is closed under non-determinism and complementation, i.e. PSPACE=NPSPACE=co-PSPACE. Since PSPACE is closed under complementation, we consider the problem of determining whether $\pi$ is not a subgame perfect equilibrium in $\mathbb{G}$.
@@
Section Goal
Section 3.2 Establish a correspondence between subgames in $\mathbb{G}$ and a set of states $R$ in $\mathbb{G}\times\pi$
Section 3.3 Establish a PSPACE algorithm for computing hitting probabilities in $\mathbb{G}\times\pi$
Section 3.4 Prove that if $\pi$ is not a subgame perfect equilibrium, this can be witnessed by a “small, local” change at a state $r\in R$ that results in a strictly higher hitting probability at $r$
Section 3.5 Compare the two hitting probabilities from Section 3.3 (no deviation) and Section 3.4 (after a “small” improvement) from the state $r\in R$ in PSPACE
Section 3.6 Consider alternate models
Figure 1. A description of the sections to follow.
@@ 3.2. Histories and Subgames
Intuitively, the idea is that this property can be determined through a reachability query that simulates the progression of an arbitrary history obtained from choosing arbitrary actions. Starting from the state $\langle v_{0},s_{0},1\rangle$, a path from $\langle v_{0},s_{0},1\rangle$ to $\langle v_{r},s_{r},n_{r}\rangle$ can be guessed state by state on-the-fly in NPSPACE=PSPACE. At each state $\langle v_{j},s_{j},n_{j}\rangle$, the algorithm can guess a successor $\langle v_{j+1},s_{j+1},n_{j+1}\rangle$ and verify in PSPACE that $\gamma(s_{j},v_{j})=s_{j+1}$, $n_{j+1}=n_{j}+1$ and there exists a global action $\theta\in\bigtimes_{k\in P(v_{j})}A_{k}$ such that $\delta(v_{j},\theta,v_{j+1})>0$, by iterating over all possibilities for $\theta$. It can also verify that $v_{j}\not\in G_{i}$. The algorithm only needs to store at most two states in its memory – the current state and the guessed successor. Since the state space $V\times S\times[F]$ is exponential in the size of the input, individual states can be encoded in PSPACE.
The set $R$ of relevant history-reachable states represents the states that occur in $\mathbb{G}\times\pi$ after some history $h$ in $\mathbb{G}$ has been observed. Therefore, if we want to analyze the payoff given to Agent $i$ in a subgame induced by a history $h$ in $\mathbb{G}$, we consider $\mathbb{G}\times\pi$ started from the state $r\in R$ that results from following the history $h$.
@@ 3.3. Computing Winning Probabilities From a State
Consider a Markov Chain $\mathcal{C}$ with two non-target states $x$ and $y$ with the following properties : The state $x$ self transitions with probability $\frac{1}{2}$ and transitions to $y$ with probability $\frac{1}{2}$. Furthermore, the state $y$ self transitions with probability $\frac{1}{2}$ and transitions to $x$ with probability $\frac{1}{2}$ – see Figure 2 for an illustration. The equations for the hitting probabilities $h_{x}$ and $h_{y}$ would be both given by the linear equation $\frac{1}{2}h_{x}+\frac{1}{2}h_{y}$, meaning that the system of equations used to solve for hitting probabilities in $\mathcal{C}$ would be underdetermined and have infinite solutions.
[…]
Figure 2. An illustration of Example $3.$ The entire figure represents the Markov Chain $\mathcal{C}$, which has two disconnected components: $\mathcal{C^{\prime}}$ and the two non-target states $x$ and $y$. We wish to calculate the hitting probabilities for some state $c\in\mathcal{C^{\prime}}$ by setting up a system of linear equations. If we set the hitting probabilities up in a naïve way, we get $h_{x}=h_{y}=\frac{1}{2}h_{x}+\frac{1}{2}h_{y}$, a pair of equations with an infinite number of solutions. Thus, the system of equations to solve for the hitting probabilities in $\mathcal{C}$ is underdetermined.
The Markov Chain $\mathbb{G}\times\pi$, however, has a special structure. Since it explicitly tracks the time index $n\in[F]$, this means that it contains no cycles – states with time index $n$ can only transition to states that have time index $n+1$.
@@
We demonstrate this by backwards induction. Consider the set of states $\langle v,s,F\rangle$, i.e. the set of states with time index $n=F$. These states have no outgoing transitions, so we have $h_{\langle v,s,F\rangle}=1$ if $v\in G_{i}$ and 0 otherwise. Therefore, all states of the form $\langle v,s,F\rangle$ have unique solutions for their hitting probabilities.
@@
Solving a system of explicitly given linear equations with a unique solution belongs to the class NC2 [Csa76], a subset of the class NC of problems solvable in parallel poly-logarithmic time using polynomially many processors [Coo85]. The standard solution concept for a system of linear equations is a vector $\bf v$} that associates a numerical value to each variable $v$ in the system of equations. Since NC2 is a complexity class containing decision problems [Joh90], the algorithm in [Csa76] technically takes as input a variable $v$ and an index $k$ and decides the value of $k$-th bit of $v$ in $\bf v$}. The class NC2 itself contained in log2-SPACE [Joh90], meaning that we can create a SPACE($\log^{2}n)$ Turing Machine $\mathcal{T}$ that is capable of simulating the NC2 algorithm presented in [Csa76]. The machine $\mathcal{T}$ takes a variable $v$, and an index $k$ as input, and it simulates the relevant processors of the NC2 algorithm to decide whether the $k$-th bit of the value of $v$ in the solution vector $\bf v$} is $1$ or 0. Since $\mathcal{T}$ runs in poly-logarithmic space, $\mathcal{T}$ would have a read-only input tape that explicitly describes the system of equations, the variable $v$, and the index $k$, alongside an additional separate work tape that uses $O(\log^{2}n)$ space [Sip06].